# TMT: Tri-Modal Translation between Speech, Image, and Text by Processing Different Modalities as Different Languages

### v1

@@ Appendix C Ablation on Speech and Image Tokens
Appendix D Example of Back Translation
We display the data obtained via BT on ImageNet and CommonVoice. For visualization purposes, we also plot the mel-spectrogram of audio. Fig. 4 and Fig. 5 show the results of BT on ImageNet dataset. By employing image-only uni-modal dataset, we can augment the paired data of text-image and speech-image using BT. The back translated dataset is utilized to further train our intermediate TMT model on Text- and Speech-driven Image Synthesis (text-to-image and speech-to-image MMT) tasks. Fig. 6 and Fig. 7 show the results on CommonVoice dataset. With the text-audio bi-modal dataset, we can build tri-modal paired dataset. The back translated dataset is used for Image Captioning and Image-to-Speech Captioning (image-to-speech and image-to-text MMT tasks) along with ASR and TTS (text-to-speech and speech-to-text MMT tasks).
Figure 4: Examples of back translated data on ImageNet dataset. The source text data is generated from the target image of ImageNet, building text-to-image pairs.
Figure 5: Examples of back translated data on ImageNet dataset. The source speech data is generated from the target image of ImageNet, building speech-to-image pairs.
Figure 6: Examples of back translated data on CommonVoice dataset. The source image data is generated from the target text of CommonVoice, building image-to-text pairs.
Figure 7: Examples of back translated data on CommonVoice dataset. The source image data is generated from the target speech of CommonVoice, building image-to-speech pairs.
Appendix E Related Work for Each MMT
E.1 Image Captioning
Image Captioning Xu et al. (2015) is a task to describe input image content in natural language. It has shown remarkable progress with the development of computer vision and natural language processing technologies. Recent methods showed that vision-language pre-training can significantly improve image captioning performances over training the model from scratch Wang et al. (2022). For example, Radford et al. (2021); Jia et al. (2021) proposed cross-modal contrastive learning, and Li et al. (2021); Kim et al. (2021b); Li et al. (2022, 2023a) proposed multi-task pre-training which contains such as image-text matching, masked language modeling, and contrastive learning. Basically, the task can be regarded as an MMT problem of image-to-text translation.
Similar yet distinct from image captioning, image-to-speech captioning Hsu et al. (2021b) aims to directly generate a spoken description of the input image. As the task does not go through with text modality, it is regarded as a challenging problem due to insufficient supervision and the complex nature of speech. Recent methods tried to mitigate the challenge by representing speech in pseudo-text format and showed promising results Hsu et al. (2021b); Effendi et al. (2021); Kim et al. (2024a). This task falls in text-to-speech translation.
E.2 Language-driven Image Synthesis
With the great development of visual codebook representation Van Den Oord et al. (2017); Esser et al. (2021) and diffusion Ho et al. (2020); Rombach et al. (2022), image generation has become more and more realistic. Especially, Ramesh et al. (2021, 2022); Saharia et al. (2022); Chang et al. (2023a) showed that generative models can now freely create images based on text descriptions by conditioning the text representations into the decoding process, such as diffusion process. They are the representatives of text-to-image translation tasks.
Different from text-based image synthesis, direct speech-to-image synthesis Li et al. (2020); Wang et al. (2020) has not well been addressed. Since many languages have no writing systems Lee et al. (2022b), text-based systems cannot be applied to such languages. Hence, the demand for technology enabling spoken language-driven image synthesis, the speech-to-image translation, is expected to increase.
Appendix F Details for Data Size (bits) Comparison
Table 1 shows the data size comparisons in bits between different data types of each modality. By employing speech tokens rather than directly using raw audio or its Mel-spectrogram, we significantly reduce the data size to less than 0.2% compared to other formats. Consequently, 1-second speech tokens contain a similar number of bits as text comprising approximately 50 characters. Furthermore, by utilizing image tokens instead of raw images, only 0.035% of the bits are needed to represent the original image content. This substantial reduction equates one image to the text length of 52 characters. Therefore, by representing and employing the tokenized data instead of directly utilizing their raw format, we can efficiently train our TMT system like training a text-only system with reduced data storage requirements and computing resources.
Appendix G Full Results of Back Translation
  Task Method BLEU-4 METEOR ROUGE CIDEr SPICE
Im. normal-$\rightarrow$ Txt. TMT w/o BT 33.9 29.3 56.7 116.7 22.3
TMT 34.3 29.2 56.9 117.7 22.3
Im. normal-$\rightarrow$ Sp. TMT w/o BT 25.0 22.8 48.9 79.5 16.6
TMT 25.8 23.3 49.5 81.3 16.8
Table 11: Performance comparisons of with BT and without BT on Image captioning and Image-to-speech captioning performances on COCO.
Table 11 shows the performances of image captioning and image-to-speech captioning on COCO in all 5 metrics when the BT is applied and not. By employing pseudo multi-modal pair datasets obtained through BT, we can consistently improve the performances overall. Please note that we can produce an unlimited number of pseudo-pair data through the multi-modal BT which may improve the performance more.

### v2

@@ IV-D3 Ablation on Speech and Image Tokens
IV-D4 Ablation Study on Multi-modal Translation Module
We perform ablation study by using different models for the multi-modal encoder-decoder module. Firstly, we investigate different initializations of translation module using BART [51], mBART [72], and BERT [50]. Specifically, the models are initialized using one of the pre-trained models and trained on 6 tasks for 300k steps on COCO and Flickr8k dataset. The comparison results on image captioning and image-to-speech captioning are shown in Table X. Both BART and BERT initialized models achieve high performances on two captioning tasks, while multilingual pre-trained model, mBART achieves inferior performances on both tasks. This might be contributed to the fact that COCO and Flickr8k are English only databases. Considering the performance gap between BERT and BART initialized models in image captioning task is much bigger than that of image-to-speech captioning, we use BERT for our encoder-decoder initialization.
Furthermore, we analyze the effects of using different model sizes for the translation module. To this end, we compare four different types of translation modules of 1) BERT-base for both encoder-decoder, 2) BERT-base for encoder and BERT-large for decoder, 3) BERT-large for encoder and BERT-base for decoder, and 4) BERT-large for both encoder-decoder. BERT-base has 12 layers with 110M parameters, while BERT-large has 24 layers with 340M parameters. All models are trained on full training set including back translation. The comparison results can be found in Table XI. Interestingly, different models achieve the best performance on different tasks. For example, model with BERT-large for both encoder-decoder (i.e., Large-Large) achieves the best result on image-to-speech captioning task while showing inferior image synthesis performances. The model with BERT-large for encoder and BERT-base for decoder (i.e., Large-Base) achieves the best image synthesis performances while showing inferior image-to-speech captioning task. Scaling up the model size shows a tendency of performance improvement when we compare the performance with the BERT-base model for both encoder-decoder (i.e., Base-Base). Given that the number of parameters of Base-Base model is much smaller than others (i.e., Base-Base: 270M parameters, Large-Large: 800M parameters) and still showing comparable performances on all 6 tasks, we report its performance for other experiments.
TABLE X: Captioning tasks’ performance comparisons using different encoder-decoder initializations on COCO.
  Translation Module Image Captioning
BLEU-4 METEOR ROUGE CIDEr SPICE
BART [51] 28.9 26.6 53.2 103.2 20.4
mBART [72] 28.4 26.3 53.1 100.9 19.9
BERT [50] 31.4 28.1 54.9 108.7 21.3
  Translation Module Image-to-Speech Captioning
BLEU-4 METEOR ROUGE CIDEr SPICE
BART [51] 24.9 23.4 49.9 80.8 17.2
mBART [72] 22.6 22.6 48.1 71.8 15.9
BERT [50] 24.7 23.2 48.9 78.7 16.7
TABLE XI: Ablation study using different network configurations for multi-modal encoder-decoder on COCO and Flickr8k dataset. ‘Base’ and ‘Large’ represents BERT-Base and BERT-Large configurations, respectively. ROUGE is reported for captioning tasks, WER for ASR and TTS, and the CLIP score for image synthesis tasks.
  Models (Encoder-Decoder) COCO
[…]
Base-Base 56.9 49.5 5.15 69.57 68.88 9.54
Base-Large 57.3 49.0 4.88 69.46 68.82 10.55
Large-Base 57.0 47.7 4.96 70.49 69.74 10.21
Large-Large 57.1 50.1 5.01 69.14 68.33 10.14
  Models (Encoder-Decoder) Flickr8k
[…]
Base-Base 54.8 48.6 4.60 70.24 68.98 8.46
Base-Large 55.2 48.9 4.26 69.90 68.78 8.91
Large-Base 55.1 48.3 4.32 71.36 70.01 8.89
Large-Large 54.9 49.8 4.42 69.73 67.94 8.90
TABLE XII: Effectiveness of Back Translation (BT) on COCO. ROUGE is reported for captioning tasks.
[…]
TMT w/o BT 56.7 48.9 5.24
TMT 56.9 49.5 5.15
[…]
TMT w/o BT 69.48 68.67 9.78
TMT 69.57 68.88 9.54
TABLE XIII: Human evaluation score (MOS) comparisons with 95% confidence interval, and Neural MOS scores in image-to-speech captioning and TTS on COCO.
  Methods MOS Neural MOS
Natural Descriptive MOSNet $\uparrow$ SpeechLM $\downarrow$
$\bullet$ Image-to-Speech Captioning
Hsu et al[34] 2.70±0.38 2.96±0.22 2.51 4.23
Kim et al[13] 3.71±0.25 3.41±0.23 4.30 4.16
[…]
$\bullet$ Text-to-Speech Synthesis (TTS)
VITS [84] 3.85±0.25 - 4.64 4.02
[…]
IV-D5 Effectiveness of Back Translation
Table XII shows the result of training TMT with and without BT. By employing additional 4M back translated data, we observe consistent improvement across tasks of TMT. This shows the potential for further improvement, since the amount of additional data BT can produce is limitless.
TABLE XIV: Ablation study on using different modalities with different MMT tasks. $S,I,T$ represent Speech, Image, and Text modalities, respectively, and arrow represents the translation directions. Best is bolded and second-best is underlined.
  Task Training Methods COCO Flickr8k
[…]
I $\rightarrow$ T I $\rightarrow$ T 53.6 - - 49.7 - -
I $\rightarrow$ T + I $\rightarrow$ S 54.1 - - 50.3 - -
I $\rightarrow$ T + S $\rightarrow$ T 54.1 - - 51.1 - -
I $\rightarrow$ T + S $\rightarrow$ I 52.0 - - 46.1 - -
TMT (all 6 tasks) 54.9 - - 52.5 - -
I $\rightarrow$ S I $\rightarrow$ S 47.1 - - 45.6 - -
I $\rightarrow$ S + I $\rightarrow$ T 49.3 - - 47.7 - -
I $\rightarrow$ S + T $\rightarrow$ S 42.1 - - 40.2 - -
I $\rightarrow$ S + T $\rightarrow$ I 46.6 - - 45.3 - -
TMT (all 6 tasks) 48.9 - - 48.5 - -
T $\rightarrow$ I T $\rightarrow$ I - 68.8 - - 69.3 -
T $\rightarrow$ I + T $\rightarrow$ S - 68.6 - - 69.3 -
T $\rightarrow$ I + S $\rightarrow$ I - 68.3 - - 68.2 -
T $\rightarrow$ I + S $\rightarrow$ T - 69.0 - - 69.3 -
TMT (all 6 tasks) - 68.2 - - 68.1 -
S $\rightarrow$ I S $\rightarrow$ I - 57.5 - - 58.5 -
S $\rightarrow$ I + S $\rightarrow$ T - 67.1 - - 66.6 -
S $\rightarrow$ I + T $\rightarrow$ I - 67.2 - - 66.5 -
S $\rightarrow$ I + T $\rightarrow$ S - 49.0 - - 49.6 -
TMT (all 6 tasks) - 67.3 - - 67.0 -
S $\rightarrow$ T S $\rightarrow$ T - - 6.9 - - 6.8
S $\rightarrow$ T + S $\rightarrow$ I - - 6.2 - - 5.7
S $\rightarrow$ T + I $\rightarrow$ T - - 6.6 - - 6.3
S $\rightarrow$ T + I $\rightarrow$ S - - 6.2 - - 5.7
TMT (all 6 tasks) - - 6.5 - - 6.1
T $\rightarrow$ S T $\rightarrow$ S - - 16.0 - - 9.7
T $\rightarrow$ S + T $\rightarrow$ I - - 14.9 - - 9.2
T $\rightarrow$ S + I $\rightarrow$ S - - 11.1 - - 11.3
T $\rightarrow$ S + I $\rightarrow$ T - - 12.5 - - 9.2
TMT (all 6 tasks) - - 11.8 - - 9.5
IV-D6 Mean Opinion Score Comparisons
In order to assess the quality of generated speech in both image-to-speech captioning and text-to-speech tasks, we conduct human subjective study and neural MOS test using MOSNet [104] and SpeechLMScore [105]. Specifically, we enlisted 21 participants who are volunteers mostly in academia to evaluate the naturalness of the generated speech for both tasks, and for the image-to-speech captioning task, we also assessed descriptiveness, gauging the extent to which the speech accurately describes the input images. Participants were asked to evaluate each sample using a score between 1 and 5, with increments of 0.5. Following the previous works [34, 13], 20 samples are used for the image-to-speech captioning task. For TTS, 20 samples and 100 samples are employed for human subjective study and neural MOS, respectively. The results are shown in Table XIII. TMT achieves the best score for both human study and neural MOS scores in image-to-speech captioning. Moreover, we achieve comparable performances with VITS on TTS task. Please note that the TTS performance of TMT can’t help relying on the performance of the token-based vocoder (i.e., HiFi-GAN [96]). As VITS showed better performances than HiFi-GAN-based TTS model in general [84], this tendency could be reflected in our experiment, where TMT shows slightly worse MOS scores compared to VITS. This means that there is the possibility of the improvement in terms of naturalness of the generated speech of TMT by improving the speech token-based vocoder. Yet, through MOS evaluation, we can confirm the effectiveness of TMT in generating intelligible and natural speech.
IV-D7 Analysis on the impact of using different modalities
[…]
V Conclusion
We introduced TMT, a simple yet efficient and effective tri-modal translation model between speech, image, and text modalities. We discretized different modalities and treat them as text. Our experiments demonstrated that the three modalities could be successfully translated using a unified multi-modal encoder-decoder architecture, incorporating six MMT tasks in a single model. TMT outperformed single MMT model counterparts. Remarkably, we showed the computational efficiency of TMT by reducing the training data size (in bits) to about 0.2% for speech and 0.04% for image data.